which is it symmetric with: x-axis, y-axis or origin?
y^2=x^2+4

Answer by stanbon(75887)

which is it symmetric with: x-axis, y-axis or origin?
y^2=x^2+4
----
Rearrange:
y^2 - x^2 = 4
------------
y^2/4 -x^2/4 = 1
-----
It's an ellipse with center at (0,0) so its symmetric to all three.
---
Interchange x with -x and you get no change: so, symmetric to y-axis
Interchange y with -y and you get no change: so, symmetric to x-axis
Interchange both and you get no change; so, symmetric to origin
==================
Cheers,
Stan H.
==================